Please answer #3!

You will be using 25.00 mL of 3.00% by mass soluti
Answer #1

For the reaction

aA + bB \rightarrow cC + dD

Rate of Reaction = -1/a d[A]/dt = -1/b d[B]/dt = 1/c d[C]/dt = 1/d d[D]/dt

The rate of the reaction will be depend upon the stichiometric coefficients of the balanced reaction, this is the reason rate of change of H2O2 is twice the rate of change of O2 in the reaction

Since according to the answer 5.6 * 10^(-6) = 2 * 2.8 * 10^(-6)